Terry’s utility function over leisure (L) and other goods (Y ) is U(L, Y ) = Y + LY. The associated marginal utilities are MUY = 1 + L and MUL = Y. He purchases other goods at a price of $1, out of the income he earns from working. Show that, no matter what Terry’s wage rate, the optimal number of hours of leisure that he consumes is always the same. 


1
Expert's answer
2021-05-04T07:38:16-0400

If Terry's hourly wage rate is w, his total earnings are "(24-L)w". "PY=1", so "Y=(24-L)w" is the number of units of other products he buys.

Terry's "MRS_L,_Y" must now equal the price ratio "\frac{w}{PY}=w" at the optimal bundle. From the tangency condition we know that "\frac{Y}{1+L}=w" As a result of the tangency condition, we can conclude that both of these conditions imply. This means that "L=11.5\%" is the ideal quantity of leisure. As you can see, this is independent of the wage rate.